OpenStudy (anonymous):

Which of the following ordered pairs is a solution to the inequality y less than or greater to -2x + 10? (0, -4) (1, 13) (6, -1) (3, 8)

OpenStudy (yanasidlinskiy):

Have you tried to solve the inequality?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

@YanaSidlinskiy

OpenStudy (yanasidlinskiy):



OpenStudy (anonymous):

D?

OpenStudy (anonymous):

Hold on.. let me check to find the answer.

OpenStudy (anonymous):

Are you sure you typed the equation properly?... because none of those are the answer to y >= -2x + 10

myininaya (myininaya):

"y less than or greater to -2x + 10?" y>-2x+10 y<-2x+10 -2x+10<y<-2x+10 This can't be true unless you mean -2x+10<=y=<-2x+10 which actually implies -2x+10=y

OpenStudy (whpalmer4):

I believe the inequality is \[y \le -2x+10\]which is satisfied by only 1 of the answer choices. I think it's most likely that the OP got the initial part of the name of the inequality sign correct, as \(y \ge -2x + 10\) is satisfied by more than one answer choice.

OpenStudy (anonymous):

sorry everybody i did have a incorrect inequality